2.1 kJ/mol. The following reaction has a value of AGⓇ = CH3Br + H2S CH,SH HBr a) Calculate key at room temperature (25°C) for the value on top is -2.1kj/mol
Answer #1

a Go= -2.1 kJ molt 160 = -2°303 RT (lug key) Reg- equillebrium constant RT = 2.48 KJ molt - -2.1 KJ Imol: -2.303 X 2.48 kJ mo= 2.33 =) keq = xxx (1-2)(1-) + 1x 12 = 2.33 Taking square root on both side we get X = 11526 x= 0.604 M [CH3 Bo] = (1 - x0)